a. Increasing or decreasing or neither or both increasing and decreasing. b. Use change in input & the change in output c
Answer #1

Solucion 1234 458 -4 -4 Express the equalón y=mctb form rate of Change = change ofy - Ag change of x D x ist case rage of chaDATE Sub y=-4 in the equahon y= b - 4-6 So the required equation y = mx + b y = ox-4 - 4 - lineax_qraph - -412- 720 3628 248and case rali of change ay 20 L-10 3rd case rate of chango Ay - 28 - 14 not linear Increasing function which means the functi